OpenStudy (anonymous):

The point P is on the unit circle. Find P(x, y) from the given information. The y-coordinate of P is -1/9, and the x-coordinate is positive. P(x,y)=( ) what is the x cordinate? thank you

OpenStudy (anonymous):

Nice question, a graph would be a good beginning

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

1 = x^2 + 1/81 1 - 1/81 = x^2 80/81 = x^2 x = sqrt( 80/81) x = (4/9) sqrt(5)

OpenStudy (anonymous):

It was a pleasure to help you